OpenStudy (anonymous):

X^3 - xy + y^3= 1 Solve for dy/dx

OpenStudy (unklerhaukus):

take the derivative term by term, what is the derivative (with respect to x) of x^3

OpenStudy (unklerhaukus):

d/dx (x^3) = ?

OpenStudy (unklerhaukus):

[for the second two terms you will need to apply the product rule]

OpenStudy (anonymous):

i have 3x^2 - x (dy/dx) + 3y^2 dy/dx = 0

OpenStudy (unklerhaukus):

youv'e missed a term, d/dx (xy) =x d/dx (y) + y d/dx (x)

OpenStudy (anonymous):

ok back up is it 3x^2 - (x dy/dx + y + 1) = 3y^2 dy/dx =0

OpenStudy (anonymous):

3x^2 - (x dy/dx + y dy/dx x ) + 3y^2 dy/dx =0

OpenStudy (anonymous):

am I on the right track

OpenStudy (unklerhaukus):

* 3x^2 - (x dy/dx + y d/dx (x) ) + 3y^2 dy/dx =0

OpenStudy (unklerhaukus):

that term y d/dx (x) = y dx/dx = y (1) = y

OpenStudy (unklerhaukus):

the other terms you have done correctly

OpenStudy (unklerhaukus):

can you see what i mean?

OpenStudy (anonymous):

yea but where do I go next

OpenStudy (unklerhaukus):

3x^2 - (x dy/dx + y) + 3y^2 dy/dx =0 now you just have to solve the equation for dy/dx by rearranging

OpenStudy (unklerhaukus):

you should end up with dy/dx = some fraction

OpenStudy (anonymous):

i am really confused on the algebra piece of it

OpenStudy (unklerhaukus):

well expand the bracket, (bring the minus across those two terms)

OpenStudy (unklerhaukus):

tell me what you get after this step

OpenStudy (anonymous):

3x

OpenStudy (anonymous):

3x^2 - x dy/dx -y +3y^2dy/dx

OpenStudy (anonymous):

sorry didn't realize I forgot to paste

OpenStudy (unklerhaukus):

dont forget the equals sign

OpenStudy (anonymous):

of course

OpenStudy (unklerhaukus):

3x^2 - x dy/dx -y +3y^2dy/dx = 0 now think of the dy/dx as just a variable factor the dy/dx from two terms

OpenStudy (anonymous):

^

OpenStudy (anonymous):

oops 3x^2 - y dy/dx ( x + 3y^2) =0

OpenStudy (unklerhaukus):

not quite

OpenStudy (anonymous):

dy/dx = (-3x^2 + y) / (x + 3y^2) = 0

OpenStudy (anonymous):

dy/dx = x/y

OpenStudy (unklerhaukus):

hmm i'm not sure how you got that but its good, except this time it dosent equal zero

OpenStudy (anonymous):

oops didn't mean to put that lol

OpenStudy (unklerhaukus):

dy/dx cannot be simplified to x/y, because d is not a number that can cancel, it is an operator

OpenStudy (anonymous):

ugh what am i doing wrong

OpenStudy (anonymous):

did I have it right before I tried to simply it

OpenStudy (unklerhaukus):

3x^2 - x dy/dx -y +3y^2dy/dx = 0 3x^2-y + dy/dx (-x + 3y^2) = 0 dy/dx (-x+3y^2) = y-3x^2 dy/dx = (y-3x^2) / (-x + 3y^2) = (3x^2 -y ) / (x-3y^2)

OpenStudy (anonymous):

thanks I don't know what I did wrong